OpenStudy (anonymous):

A rectangle with a length of x - 4 and a width of 8 has a perimeter of 34 What is the value of X?

OpenStudy (lalaly):

perimeter= 2*length + 2* width 34 = 2(x-4)+2*8 34 = 2x-8+16 34=2x+8 26=2x x=13